OpenStudy (anonymous):

Is the solution to the radical equation the square root of negative 6 x equals x x = 0, x = -6 or both?

OpenStudy (anonymous):

Both. sqrt(-6x) = x (-6x) = x^2 x^2 + 6x = 0 (x)(x+6) = 0 thus x = 0, x+6 = 0 *splitting the ()() up*